OpenStudy (anonymous):

find the x and y intercepts and graph the equation plotting the intercepts. -4x+5y=-20 please show work

OpenStudy (anonymous):

to find the x intercept let y = 0 -4x + 5(0) = -20 -4x = -20 x = 5 therefore x-intercept is (5,0)

OpenStudy (anonymous):

to find the y-intercept you let x become 0 -4(0) + 5y = -20 5y = -20 can you solve for y now?

OpenStudy (anonymous):

would it be (0,-20)

OpenStudy (anonymous):

sorry i mean(-20,0)

OpenStudy (anonymous):

y would be -20/5 what is -20/5?

OpenStudy (anonymous):

ok so -4x+5y=-20 5y=4x-20 y=4/5x-4 y intercept is -4 x intercept is when y is 0 so. 0=4/5x-4 4=4/5x 5=x 5 is x intercept

OpenStudy (anonymous):

ok now im more confused bc i have two different answers

OpenStudy (anonymous):

what are you asking for x or y intercept?

OpenStudy (anonymous):

the question asks for both x and y

OpenStudy (anonymous):

x is 5 and y is -4

OpenStudy (anonymous):

how did you come up with that? i have to ge able to make a line graph